### v5

@@ 2.2 Mixed Cooperation and Noncooperation
The $1\times n$ row vector $\pi$, referred to as the authority distribution in Hu and Shapley (2003), also solves the following counterbalance equilibrium:
@@
subject to the normalization conditions $\sum\limits_{i=1}^{n}\pi_{i}=1$ and non-negativity condition $\pi_{i}\geq 0$. Numerous counterbalanced systems exist in both the physical and human worlds, such as ecological systems, the USA’s government system, and China’s five-element theory.
The counterbalance in Eq. (3) implies $\pi$’s mixed cooperative and noncooperative properties. This is a distinctive feature between comparative and competitive advantages; while both promote collaborations, competitive advantage may bring trade wars and deglobalization. For instance, between China and the USA from 1980 to 2020, there was more cooperation in the first thirty years than in the last ten years. Although unfamiliar to empirical macroeconomics, this mixture could have an edge in the future (e.g., Allen, 2000, page 147). Econometric analysis often struggles to capture the complexity of antagonism, where any estimated coefficient or implicated effect cannot be both significantly positive and significantly negative.
From a non-cooperative perspective, as $\sum\limits_{i=1}^{n}\pi_{i}=1$, an increase in $\pi_{j}$ may imply a decrease in $\pi_{i}$. Thus, theoretically, there are $n(n-1)/2$ potential trade conflicts, either small or large, within the trade system. Most of these are minor trade disagreements or disputes, making the “war” metaphor hyperbolic in these cases. However, it is worth mitigating even minor disputes before they become problematic or catastrophic. Additionally, the increase in $\pi_{j}$ is not evenly redeemed by other economies; they often wax and wane with some common causality initiated by changes in the matrix $P$. In Sections 3 and 5, we distinguish the first move off the diagonal and that on the diagonal of $P$, respectively, for bilateral trade wars and globalization.
From a cooperative perspective, country $i$ should assist country $j$ in improving $\pi_{j}$ whenever $P_{ji}>0$, as $\pi_{j}P_{ji}$ is part of $\pi_{i}$, as shown in:
@@
from Eq. (3). Eq. (3) also implies that $\pi=\pi P=\pi P^{2}=\cdots=\pi P^{t},$ where
@@
for any $t=1,2,3,\cdots.$ When $\pi_{i}>0$, as $\lim\limits_{t\to\infty}\left(P^{t}\right)_{ji}=\pi_{i}>0$, therefore, $\left(P^{t}\right)_{ji}>0$ for all large $t$ and all $j\in\mathcal{N}$. Thus, Eq. (5) further implies that country $i$ should assist all other countries in the trade system, including the poorest and least competitive ones, to enhance their $\pi_{j}$. Hence, the double-edged counterbalance suggests that a carrot-and-stick approach would better leverage country $i$’s position in the trade system when dealing with its trade partner $j$. The trade-off is determining how much it should assist or contest $j$ without sacrificing its national interests and potentially increasing $\pi_{i}$.
@@
Figure I: Dynamics of authority flow for country $i$ and $j$.
We can conceptualize $\pi_{i}$ as a container or pool within the dynamics of counterbalance equilibrium, where comparative and competitive advantages serve as the primary driving forces for the flows. In the inflow equation Eq. (4), country $i$ accumulates or absorbs authority from other countries through its direct influences. It derives more authority from influential trade partners (i.e., $\pi_{j}$ is large) than from non-influential ones, all else being equal. Additionally, it derives more authority from countries on which it has a significant direct influence (i.e., $P_{ji}$ is large), all else being equal. In Figure I, country $i$ imports goods from $j$ through the $i$th column of $P$, and thus, authority flows from $j$ to $i$ along the same column.
Conversely, country $i$ also contributes to other countries, as illustrated in the following outflow equation:
@@
The larger $P_{ij}$ or $\pi_{i}$, the more $i$ contributes to $j$’s authority $\pi_{j}$, all else being equal. In Figure I, country $i$ redistributes $\pi_{i}$ through the $i$th row of $P$, according to its shares of exports.
In the accounting of $\pi$, the two countries reach a break-even point when:
$\pi_{i}P_{ij}=\pi_{j}P_{ji}.$ (6)
Intuitively, if $\pi_{i}P_{ij}<\pi_{j}P_{ji}$ or $\pi_{i}P_{ij}>\pi_{j}P_{ji}$, then country $i$ either takes more authority from or gives more to country $j$, respectively. For country $i$, a straightforward implication of Eq. (6) is to decrease or increase its imports from $j$ if $\pi_{i}P_{ij}<\pi_{j}P_{ji}$ or $\pi_{i}P_{ij}>\pi_{j}P_{ji}$, respectively. It can also adjust its exports to $j$ accordingly. However, the outflow from $i$ to $j$ does not necessarily match the inflow. With the involvements of third parties, the inflows and outflows meet at the equilibrium Eq. (3) for $i$, i.e. $\pi_{i}P_{ij}+\sum\limits_{k\neq j}\pi_{i}P_{ik}=\pi_{j}P_{ji}+\sum\limits_{k\neq j}\pi_{k}P_{ki}=\pi_{i}$. For the same reason, it could be overly restrictive for all countries to balance their bilateral trades. For example, country $i$ can still maintain an overall zero net balance when it has a trade surplus with $j$ and a trade deficit with $k$ of the same amount.
@@
Finally, Eq. (3) implies that $\pi$ is the unit-sum row eigenvector of the largest row eigenvalue of $P$. It represents a type of eigenvector centrality, a subclass of network centrality. In the literature, eigenvector centrality is the dominant eigenvector for an adjacency or relation matrix, which often has a zero diagonal (e.g., Bonacich, 1987). Our $P$ has a nonzero diagonal, allowing domestic and foreign goods to compete in the domestic market, and the eigenvector $\pi$ measures competitiveness (see Section 2.3). The same domestic products are used both domestically or overseas. Hu (2020) also retains a nonzero diagonal when any accepted student selects only one college to attend from multiple offers; thus, the selected college on the diagonal competes with the others that also accept the student and are off the diagonal. Furthermore, $P$ normalizes the scales of economies such that each row has a unit sum. Consequently, $P$ has the largest row eigenvalue $1$, common to all stochastic matrices.
2.3 $\pi$ as Competitive Advantages
@@
The competitive advantage of country $i$ refers to its ability to outperform its competitors within the trade system. For any product exported from country $i$ to country $j$, the competitors of country $i$ include all other countries. Thus, its competitive advantage extends from bilateral country relations to a relationship with the entire system — it is the ability to surpass all other countries in the $j$th market for any $j\in\mathcal{N}$. In contrast, comparative advantage is an economy’s leverage of specialization, resource endowment, and technology over the trade partner’s opportunity cost; it is a bilateral relationship. Therefore, globalization may have made competitive advantages more prevalent than comparative advantages.
@@
According to the literature (e.g., Porter, 1985; Stutz and Warf, 2010), competitive advantages rest on the direct control of competitive factors of production, such as access to rare resources or advanced technology, cheap but highly skilled labor, land, market affordability, and low cost of capital. These factors, imported from outside or manufactured domestically, are assembled in the final products. Thus, the imported goods carry the original producer’s competitiveness and should be attributed to the original maker. In the smartphone example, the competitive factors include the chips and patents from the USA, the labor from China, the precision manufacturing in Japan, and the sophisticated cameras from Korea. Various studies have also described qualitative strategies to enhance firm competitiveness. For example, Porter (1985) lists three generic strategies: cost leadership, differentiation, and focus.
We find that $\pi$ quantifies national competitiveness. To introduce the quantitative competitiveness, we ignore the home bias in trade (cf., McCallum, 1995) and make a strong assumption for the time being: no bilateral trade deficit nor surplus. Denote $\tau_{j}$ as the $j$th row of $P$. Because of no bilateral surplus nor deficit, $\tau_{j}$ except for its $j$th element equals the imports from other countries to $j$, all in fractions of $j$’s production. So, it measures other countries’ aggregate comparative advantages over country $j$. Besides, all other countries also compete for the comparative advantages over $j$; a large value in $\tau_{j}$ means solid competitiveness to occupy the $j$th market and win a lion’s share of $j$’s imports. If we use the $j$th market as the reference to rank the competitiveness of all countries, then the competitiveness score is the vector $\tau_{j}$, which has a unit sum. Yet, the $j$th value is not excluded from $\tau_{j}$, because domestic products also compete with country $j$’s imports when consumers make their consumption choices, and they have no home bias.
@@
We apply endogenous weighting to the references to combine these scores $\tau_{1},\tau_{2},\cdots,\tau_{n}$. These individual cardinal utilities are not comparable because it is more challenging to occupy a competitive market than a less competitive one. We make them comparable by weighting, and the weights are not exogenously assigned but endogenously determined. More weight should be placed on a more competitive reference market than on a less competitive one. Let the unknown weights be $\rho=(\rho_{1},\rho_{2},\cdots,\rho_{n})$ with $\sum\limits_{j=1}^{n}\rho_{j}=1$ and $\rho_{j}\geq 0$ for all $j\in\mathcal{N}$. Then, we have a weighted competitiveness score $\sum\limits_{j=1}^{n}\rho_{j}\tau_{j}=\rho P$. Both the weighted score and the weight vector quantify the competitiveness of the countries. To be consistent, the weighted average $\rho P$ should be a positive multiple of the weight vector $\rho$. By Theorem 1, $\rho$ is just $\pi$.
@@
If $\rho P=c\rho$ for some constant $c>0$, then $c=1$ and, therefore, $\rho=\pi$.
@@
In the endogenous weighting, worldwide consumers unconsciously participate in evaluating national competitiveness, encompassing all products and services. Essentially, countries as a whole unknowingly assess themselves without external interventions from business interests, mainstream media, advertising, governments, or even military threats. The structure of the global trade network is largely determined endogenously by the distribution of firm productivity within individual countries. Nevertheless, government policies can influence this distribution without significantly affecting the welfare of international trade by shifting imports or exports between countries. Policies can also enhance firm productivity through tax exemptions, subsidies, and imposing heavy tariffs on foreign products. However, national competitiveness $\pi_{i}$ is not merely the sum of the individual competitiveness of domestic firms; firm competitiveness are weighted in $\pi$ by consumers, wealthy or poor, domestic or overseas. Affordability thus plays a vital role in $\pi$. The positive weights imply a positive correlation between firm and national competitiveness.
Additionally, $\pi$ represents a von Neumann-Morgenstern Bergson social welfare function, as a positively weighted sum of individual von Neumann-Morgenstern utility functions $\tau_{1},\tau_{2},\cdots,\tau_{n}$. Thus, the weighted utility function $\pi$ is determined not by the consumption of production factors or the added value, but by consumer preferences, judged independently of national interests or preferences. Pareto optimality is achieved when weighting these individual utilities: $\pi_{j}\geq\pi_{k}$ if $P_{ij}\geq P_{ik}$ for all $i\in\mathcal{N}$. The optimality remains invariant if any individual utility function is scaled by a positive constant or added by a constant, such as scaling $\tau_{i}$ by $g_{i}$. Harsanyi (1955) and Hammond (1987) provide additional ethical, risk, and utilitarianism connotations for this type of social welfare function. With the weights being $\pi$ itself, $\pi$ exhibits some special properties, discussed in Section 2.4.
However, the social welfare function $\pi$ significantly departs from widely-used utility or welfare functions in the literature on trade wars or globalization (e.g., Grossman and Helpman, 1995; Harrison and Rutstrom, 1991; Ossa, 2014). It is not explicitly expressed in a specific functional form of production factors and total factor productivity but implied from the counterbalance equilibrium, as indirect imports also impact the final products or services. Exported goods and services have different utilities for exporting and importing countries. Therefore, each country has a unique utility function, and a single utility function cannot capture the diversity of all countries. We use fewer assumptions to accomplish $\pi$ by ignoring how efficiently each country packages production factors into final products and how differently it trades the products with others. It is precisely derived from the data and uses no parametric models, thus no unknown parameters need to be estimated, and no residuals exist to differentiate the data from a parametric model. Additionally, according to Barney and Felin (2013), micro-foundations research is still heavily debated, with management, strategy, and organization scholars having varying views on the link between macroeconomic phenomena and microeconomic agents’ behaviors. Due to broad firm and national heterogeneity, micro-founded strategic movements are particularly challenging on the network with flows of millions of products over tens of thousands of directional edges. With this complexity, a certain level of abstraction and ambiguity is necessary to grasp the essential parts of the network.
@@
The index $\pi_{i}$ contains a unique component, $\pi_{i}P_{ii}$, which does not originate from external source, as indicated by Eq. (4). This encompasses, but is not limited to, raw oil fields, natural gas reserves, rare minerals, and patented technologies, all of which remain integral to the nation’s core competitive factors. A global value chain typically involves finite steps but also includes indefinite domestic loops when tracing all the added values of the final products. For instance, in the smartphone example, the education, skills, and experience of manufacturing labor are ultimately developed domestically over many years, contributing to the final product. Additionally, numerous U.S. workers develop patents and design software domestically for extended periods. These inner loops affect other countries but not vice versa; this aspect of competitiveness never flows in from external sources. Consequently, social infrastructure, political institutions, monetary and fiscal policies, microeconomic environments, and other quality determinants that facilitate business all contribute to national competitiveness (e.g., Delgado et al., 2012).
@@
For a specific global value chain, at the other end of the unique component is the design and assembly of the final product. Product leadership bridges the market, consumers, strategic direction, technology innovation, sales, and production factors together; this leadership represents another type of competitiveness. Between the end product and the raw resources in the chain are a series of value-added processes for intermediate goods. Their competitiveness lies in cost efficiency or cost leadership, including the cost of financing. Technological advancements reduce labor costs, and public safety and social stability mitigate the risk of costs.
Eq. (2) indicates a robust positive correlation between $\pi_{i}$ and $g_{i}$, the breadth of the economy. After adjusting for the sizes of $g_{i}$ in $P$, the interaction effects between competitiveness and economies of scale remain, thereby conferring some cost advantages. Nations export goods that are abundant and import goods that are scarce domestically. With significant abundance and shortage, country $i$ could have substantial nonzero elements in the $i$th column and row of the matrix $P$, resulting in considerable nonzero items summed in Eq. (2). A large economy also generates a high differentiation of goods and services, which offer location advantages and minimal transaction costs for domestic consumers. Free from tariffs and overseas shipping, these goods also possess a price advantage. Consequently, $P_{ii}$ is often the largest in the $i$th row of $P$, irrespective of home bias. Therefore, a vast economy with a comprehensively integrated industrial system (such as China and the USA) is likely, though not necessarily, to have a large value in $\pi$. A country with a single industrial structure is susceptible to intense competition, and its $\pi_{i}$ is fragile and volatile when competing with the USA for competitiveness.
Hence, we use $\pi_{i}$ as an objective for country $i$ when formulating trade policies. It quantifies the economy’s depth and its interplay with breadth, encompassing both cooperation and non-cooperation with other nations. This measure is endogenously derived from the comparative advantages of all products and services. Given that each country has unique locations, natural and human resources, and economic systems, a single utility function that assumes one stereotype for all economies may be insufficient to capture this diversity. Our study acknowledges that each country, along with its trade stance with partners, is distinct. Each contributes inclusiveness value to the global economy, where international trade is a cornerstone.

### v6

@@ 2.2 Mixed Cooperation and Noncooperation
The $1\times n$ row vector $\pi$, referred to as the authority distribution in Hu and Shapley (2003), also satisfies the following counterbalance equilibrium:
@@
subject to the normalization conditions $\sum\limits_{i=1}^{n}\pi_{i}=1$ and non-negativity condition $\pi_{i}\geq 0$. Counterbalanced systems are prevalent in both the physical and human domains — examples include ecological systems, the U.S. government structure, and China’s five-element theory.
The counterbalance expressed in Eq. (3) reflects the mixed cooperative and noncooperative nature of $\pi$. This duality distinguishes comparative advantage from competitive advantage: while both encourage collaboration, competitive advantage can also lead to trade wars and deglobalization. For example, in the trade relationship between China and the United States from 1980 to 2020, the first thirty years were marked by greater cooperation than the last ten.
Although this mixture of cooperation and noncooperation is less familiar to empirical macroeconomics, it may offer valuable insights moving forward (e.g., Allen, 2000, page 147). Also, econometric models often struggles to capture the unity of opposites, as estimated coefficients or effects cannot simultaneously be significantly positive and significantly negative.
From a noncooperative standpoint, given that $\sum\limits_{i=1}^{n}\pi_{i}=1$, an increase in $\pi_{j}$ may imply a decrease in $\pi_{i}$. Theoretically, this leads to $\dfrac{n(n-1)}{2}$ potential trade conflicts — ranging from minor disagreements to major disputes — within the trade system. Nonetheless, even minor disputes are worth addressing before they escalate into serious or catastrophic issues. Moreover, the increase in $\pi_{j}$ is not uniformly offset by reductions in other economies. These shifts often follow shared causal patterns triggered by changes in the matrix $P$. In Sections 3 and 5, we differentiate between the first move off the diagonal and the move along the diagonal of $P$, corresponding to bilateral trade wars and globalization, respectively.
From a cooperative perspective, country $i$ should assist country $j$ in improving $\pi_{j}$ whenever $P_{ji}>0$, because $\pi_{j}P_{ji}$ contributes to $\pi_{i}$, as shown in:
@@
from Eq. (3). Eq. (3) also implies that $\pi=\pi P=\pi P^{2}=\cdots=\pi P^{t}$ where:
@@
for any $t=1,2,3,\cdots.$ When $\pi_{i}>0$, as $\lim\limits_{t\to\infty}\left(P^{t}\right)_{ji}=\pi_{i}>0$, it follows that $\left(P^{t}\right)_{ji}>0$ for all sufficiently large $t$ and for all $j\in\mathcal{N}$. Thus, Eq. (5) further implies that country $i$ should assist all other countries in the trade system — including the poorest and least competitive ones — to enhance their $\pi_{j}$.
Therefore, the double-edged counterbalance suggests that a carrot-and-stick approach would better leverage country $i$’s position in the trade system when dealing with its trade partner $j$. The key trade-off lies in determining how much assistance or contestation to provide to $j$ without sacrificing national interests while potentially increasing $\pi_{i}$.
@@
Figure I: Dynamics of authority flow for country $i$ and $j$
We can conceptualize $\pi_{i}$ as a container or pool within the dynamics of counterbalance equilibrium, where comparative and competitive advantages serve as the primary forces driving the flows. In the inflow equation Eq. (4), country $i$ accumulates or absorbs authority from other countries through its direct influences. It gains more authority from influential trade partners (i.e., when $\pi_{j}$ is large) than from less influential ones, all else being equal. Additionally, it derives more authority from countries on which it exerts significant direct influence (i.e., when $P_{ji}$ is large), again holding other factors constant. As shown in Figure I, country $i$ imports goods from $j$ through the right column, and thus authority flows from $j$ to $i$ along the same column.
Conversely, country $i$ also contributes to other countries, as illustrated in the outflow equation:
@@
The larger $P_{ij}$ or $\pi_{i}$, the more $i$ contributes to $j$’s authority $\pi_{j}$, all else being equal. In Figure I, country $i$ redistributes $\pi_{i}$ through the bottom row, according to its export shares.
These two countries reach a break-even point when:
$\pi_{i}P_{ij}=\pi_{j}P_{ji}$ (6)
where $\pi_{i}P_{ij}$ and $\pi_{j}P_{ji}$ represent the authority flows from $i$ to $j$ and from $j$ to $i$, respectively. Intuitively, if $\pi_{i}P_{ij}<\pi_{j}P_{ji}$ or $\pi_{i}P_{ij}>\pi_{j}P_{ji}$, then country $i$ either gains more authority from or gives more to country $j$, respectively. For country $i$, a straightforward implication is to decrease or increase its imports from $j$ if $\pi_{i}P_{ij}<\pi_{j}P_{ji}$ or $\pi_{i}P_{ij}>\pi_{j}P_{ji}$, respectively. It can also adjust its exports to $j$ accordingly.
However, the outflow from $i$ to $j$ does not necessarily match the inflow. With the involvement of third parties, inflows and outflows meet at the equilibrium described by Eq. (3) for $i$:
$\pi_{i}P_{ij}+\sum\limits_{k\neq j}\pi_{i}P_{ik}=\pi_{j}P_{ji}+\sum\limits_{k\neq j}\pi_{k}P_{ki}=\pi_{i}.$
For the same reason, requiring all countries to balance their bilateral trades would be overly restrictive. For example, country $i$ can still maintain an overall zero net balance if it has a trade surplus with $j$ and a trade deficit with $k$ of the same amount.
@@
Finally, Eq. (3) implies that $\pi$ is the unit-sum row eigenvector corresponding to the largest row eigenvalue of $P$. This represents a form of eigenvector centrality, a subclass of network centrality. In the literature, eigenvector centrality is the dominant eigenvector for an adjacency or relation matrix, which often has a zero diagonal (e.g., Bonacich, 1987). Our matrix $P$ has a nonzero diagonal, allowing domestic and foreign goods to compete in the domestic market and the same domestic products are used both domestically and overseas. Hu (2020) also retains a nonzero diagonal when any accepted student selects only one college to attend from multiple offers; thus, the selected college on the diagonal competes with others that also accept the student and are off the diagonal. Furthermore, $P$ normalizes the scales of economies such that each row sums to one. Consequently, $P$ has the largest row eigenvalue $1$, common to all stochastic matrices.
2.3 $\pi$ as a Measure of Competitive Advantage
@@
The competitive advantage of country $i$ refers to its ability to outperform rivals within the global trade system. For any product exported from country $i$ to country $j$, the competitors of $i$ include all other countries. Thus, competitive advantage extends from bilateral relations to encompass the entire system — it reflects the ability to surpass all other countries in market $j$ for any $j\in\mathcal{N}$. In contrast, comparative advantage is based on an economy’s specialization, resource endowment, and technology relative to its trade partner’s opportunity cost, making it inherently bilateral. Globalization, however, may have amplified the importance of competitive advantages over comparative advantages.
@@
According to the literature (e.g., Porter, 1985; Stutz and Warf, 2010), competitive advantage rests on direct control of key production factors, such as access to rare resources, advanced technology, low-cost but highly skilled labor, affordable markets, and inexpensive capital. These factors — whether imported or domestically produced — are integrated into final products. Consequently, imported goods carry the original producer’s competitiveness and should be attributed to their source. For example, in smartphone production mentioned before, competitive factors include U.S. chips and patents, Chinese labor, Japanese precision manufacturing, Korean camera technology. Various studies also outline qualitative strategies to enhance firm competitiveness. Porter (1985), for instance, identifies three generic strategies: cost leadership, differentiation, and focus.
We propose that $\pi$ quantifies national competitiveness. To introduce this quantitative measure, we temporarily ignore home bias in trade (cf., McCallum, 1995) and assume no bilateral trade surplus or deficit. Denote $\tau_{j}$ as the $j$th row of $P$. Under the assumption of no bilateral surplus or deficit, $\tau_{j}$ — excluding its $j$th element — represents imports from other countries into $j$, expressed as fractions of $j$’s production. This measures other countries’ aggregate comparative advantages over $j$.
Moreover, all countries compete for comparative advantages in market $j$. A large value in $\tau_{j}$ indicates strong competitiveness in capturing a significant share of $j$’s imports. If we use market $j$ as the reference point for ranking all countries’ competitiveness, the resulting score is the vector $\tau_{j}$, which sums to one. The $j$th element is included because domestic products also compete with imports when consumers make purchasing decisions, assuming no home bias.
@@
We apply endogenous weighting to the reference scores $\tau_{1},\tau_{2},\cdots,\tau_{n}$. These individual cardinal utilities are not directly comparable because operating in a highly competitive market is inherently more challenging than in a less competitive one. To make them comparable, we introduce weights that are endogenously determined rather than exogenously assigned. Greater weight should be given to more competitive reference markets.
Let the unknown weights be $\rho=(\rho_{1},\rho_{2},\cdots,\rho_{n})$, where $\sum\limits_{j=1}^{n}\rho_{j}=1$ and $\rho_{j}\geq 0$ for all $j\in\mathcal{N}$. The weighted competitiveness score is:
$\sum\limits_{j=1}^{n}\rho_{j}\tau_{j}=\rho P.$
Both the weighted score and the weight vector quantify national competitiveness. For consistency, the weighted average $\rho P$ should be a positive multiple of the weight vector $\rho$. By Theorem 1, $\rho$ equals $\pi$.
@@
If $\rho P=c\rho$ for some constant $c>0$, then $c=1$ and therefore $\rho=\pi$.
In this endogenous weighting framework, global consumers unconsciously participate in evaluating national competitiveness across all products and services. Essentially, countries collectively assess themselves without external influence from business interests, media, advertising, governments, or military pressure.
@@
The structure of the global trade network is largely shaped by the distribution of firm productivity within individual countries. However, government policies can influence this distribution — without significantly affecting international trade welfare — by shifting imports or exports between countries (e.g., Gopinath et al., 2024). Policies can also enhance firm productivity through tax exemptions, subsidies, and imposing tariffs on foreign products.
Importantly, national competitiveness $\pi_{i}$ is not simply the sum of domestic firms’ competitiveness. Instead, firm competitiveness is weighted in $\pi$ by consumer preferences — wealthy or poor, domestic or overseas. Affordability thus plays a critical role in determining $\pi$. Positive weights imply a positive correlation between firm-level and national competitiveness.
Moreover, $\pi$ represents a von Neumann-Morgenstern Bergson social welfare function, as it is a positively weighted sum of individual von Neumann-Morgenstern utility functions $\tau_{1},\tau_{2},\cdots,\tau_{n}$. This weighted utility function $\pi$ is determined not by factor consumption or added value, but by consumer preferences, independent of national interests. Pareto optimality is achieved because:
$\pi_{j}\geq\pi_{k}\ \mathrm{if}\ P_{ij}\geq P_{ik}\ \mathrm{for}\ \mathrm{all}\ i\in\mathcal{N}.$
This optimality remains invariant if any individual utility function is scaled or shifted by a positive constant (e.g., scaling $\tau_{i}$ by $g_{i}$). Harsanyi (1955) and Hammond (1987) provide further ethical, risk, and utilitarian interpretations of this type of social welfare function.
This social welfare function $\pi$ differs significantly from widely used utility or welfare functions in the literature on trade wars and globalization (e.g., Grossman and Helpman, 1995; Harrison and Rutstrom, 1991; Ossa, 2014). Unlike traditional approaches that specify a functional form based on production factors and total factor productivity, our formulation is implicitly derived from the counterbalance equilibrium, accounting for the indirect effects of imports on final goods and services. Because exported goods and services yield different utilities for exporting and importing countries, each country possesses a distinct utility function. Consequently, a single utility function cannot capture the heterogeneity across nations.
Our approach minimizes assumptions by disregarding how efficiently countries transform production factors into final products or how they engage in trade. The index $\pi$ is constructed directly from observed data without relying on parametric models. As a result, there are no unknown parameters to estimate and no residuals to reconcile with a model specification.
Furthermore, as Barney and Felin (2013) note, the debate over micro-foundations remain unresolved, with scholars in management, strategy, and organization offering divergent views on the relationship between macroeconomic phenomena and micro-level behaviors. Given the vast heterogeneity among firms and nations, modeling micro-founded strategic interactions within a network comprising millions of product flows across tens of thousands of directional edges is particularly challenging. A certain degree of abstraction is therefore necessary to capture the essential features of such a complex system.
The index $\pi_{i}$ includes a unique component, $\pi_{i}P_{ii}$, which originates entirely from domestic sources, as indicated in Eq. (4). This component encompasses critical resources such as oil fields, natural gas reserves, rare minerals, and patented technologies — factors that remain central to a nation’s competitive advantage.
@@
A global value chain typically consists of a finite sequence of steps but also includes indefinite domestic loops when tracing the value added to final products. For example, in the case of smartphones, the education, skills, and experience of manufacturing labor are developed domestically over many years, contributing to the final product. Similarly, U.S. workers invest significant time in developing patents and design software domestically. These internal loops influence other countries but are not reciprocally affected, underscoring the role of domestic competitiveness $\pi_{i}P_{ii}$. Elements such as social infrastructure, political institutions, monetary and fiscal policies, and the quality of the microeconomic environment all contribute to national competitiveness (e.g., Delgado et al., 2012).
@@
Within a specific global value chain, the opposite end of this unique component involves the design and assembly of the final product. Product leadership integrates market orientation, consumer preferences, strategic direction, technology innovation, sales, and production factors — representing another dimension of competitiveness. Between raw resources and the final product lies in a series of value-added processes for intermediate goods, where competitiveness is driven by cost efficiency, including financing costs. Technological progress reduces labor costs, while public safety and social stability mitigate risk-related expenses.
Eq. (2) reveals a strong positive correlation between $\pi_{i}$ and $g_{i}$, which reflect the depth and breadth, respectively, of a country’s economy. After adjusting for the magnitude of $g_{i}$ in matrix $P$, the interaction between competitiveness and economies of scale persists, offering notable cost advantages. Countries tend to export goods they possess in abundance and import those they lack. When these are significant, country $i$ will exhibit substantial nonzero elements in both the $i$th column and row of matrix $P$, contributing meaningfully to the summation in Eq. (2).
Large economies also foster a high degree of product and service differentiation, which provides location-specific advantages and reduces transaction costs for domestic consumers. These goods, being exempt from tariffs and international shipping fees, enjoy a price advantage. As a result, $P_{ii}$ is often the largest in the $i$th row of $P$, regardless of any home bias. Therefore, a large and comprehensively integrated industrial system — such as those found in China or the United States — is likely, though not guaranteed, to exhibit a high value in $\pi$. In contrast, countries with narrow industrial structures are more vulnerable to competitive pressures, and their $\pi_{i}$ tend to be fragile and volatile, especially when competing with the U.S. in terms of competitiveness.
We thus adopt $\pi_{i}$ as a strategic objective for country $i$ in the formulation of trade policies. It encapsulates both the depth and breadth of the economy, reflecting cooperative and non-cooperative dynamics with other nations. This metric is endogenously derived from the comparative advantages across all goods and services. Recognizing that each country possesses unique geographic, natural, human, and institutional characteristics, we argue that a single utility function cannot adequately represent the diversity of global economies. Our study embraces the distinctiveness of each nation and its trade relationships, emphasizing the inclusive value each contributes to the global economy — where international trade remains a foundational pillar.